OpenStudy (anonymous):

What is the mean of this set of data: 78, 79, 80, 80, 89, 89, 92, 93, 94, 95, 95, 95? Express your answer as a decimal number rounded to the nearest tenth.

OpenStudy (confluxepic):

\[\huge \color{blue}{Welcome~To~OpenStudy!!!}\]

OpenStudy (anonymous):

first add them all together

OpenStudy (confluxepic):

Then divide them by the number of numbers.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 1059/12 =88.30

OpenStudy (confluxepic):

\[\frac{ 78+79+80+80+89+89+92+93+94+95+95+95 }{ 12}\]

OpenStudy (confluxepic):

@beepbooooop

OpenStudy (anonymous):

i like your pic confluxepic

OpenStudy (confluxepic):

Thanks.

OpenStudy (confluxepic):

\[\frac{ 78+79+80+80+89+89+92+93+94+95+95+95 }{ 12}=\frac{ 1059 }{12}\]

OpenStudy (confluxepic):

@beepbooooop

OpenStudy (confluxepic):

\[\frac{ 1059 }{ 12 }=88.25\]

OpenStudy (confluxepic):

@beepbooooop

OpenStudy (aum):

Don't forget to round it to the nearest tenth.

OpenStudy (anonymous):

88.30

OpenStudy (aum):

88.3

OpenStudy (anonymous):

lol still

OpenStudy (coconutjj):

<b>Hello</b>

OpenStudy (coconutjj):

darnit html doesnt work in openstudy